How do I solve this equation for the value of x?
14/x + (7x – 5) / (x + 2) = 7

Answer by mananth(16946)

14/x + (7x – 5) / (x + 2) = 7
add the two fractions LCM = x(x+2)
14(x+2)+(7x-5)*x / x(x+2)=7
14x+28+7x^2-5x / x^2 +2x =7
14x+28+7x^2-5x= 7(x^2+2x)
14x+28+7x^2-5x= 7x^2+14x
5x=28
x=28/5